celticcat:

A stone is dropped from the top of a building and two seconds later another stone is thrown vertically downwards at a speed of 25 m/s. Both stones reach the ground at the same time. Find the height of the building

jhonyy9:

try drawn this cases

jhonyy9:

celticcat are you here ?

jhonyy9:

how you think what is the speed of first stone what not was thrown vertically downwards ?

jhonyy9:

 how you think what mean ,how can you writing that the second stone was thrown 2 second later ?

jhonyy9:

you can check it here : https://en.wikipedia.org/wiki/Velocity

celticcat:

Sorry I got called away. I used the following equations of motion. General form h(t) = ut + 1/2gt^2 where h = height, u = initial velocity and g = acceleration due to gravity, t = time in seconds. Using the given values we have h(t) = 1/2 * 9.8 t^2 h = 25t + 1/2*9.8 * (t - 2)^2. If these are correct then t comes out negative which makes me think that the values in the question are invalid. It'a question in a textbook.